OpenStudy (anonymous):

Solve and check 29x^2-20x=0

OpenStudy (anonymous):

pull an x out

OpenStudy (anonymous):

would you set them both to zero, such as 29x^2=0 and 20x=0?

OpenStudy (anonymous):

absolutely not

OpenStudy (anonymous):

okay then how would you pull an x out?

OpenStudy (anonymous):

x(29x-20)=0

OpenStudy (anonymous):

x=0 x=20/29

OpenStudy (anonymous):

can you show me how your worked it out, I understand how you got the answer but the work behind it?

OpenStudy (anonymous):

like a proof?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

or the step you took to get the solution

OpenStudy (nathan917):

Simplifying 29x2 + -20x = 0 Reorder the terms: -20x + 29x2 = 0 Solving -20x + 29x2 = 0 Solving for variable 'x'. Factor out the Greatest Common Factor (GCF), 'x'. x(-20 + 29x) = 0 Subproblem 1 Set the factor 'x' equal to zero and attempt to solve: Simplifying x = 0 Solving x = 0 Move all terms containing x to the left, all other terms to the right. Simplifying x = 0 Subproblem 2 Set the factor '(-20 + 29x)' equal to zero and attempt to solve: Simplifying -20 + 29x = 0 Solving -20 + 29x = 0 Move all terms containing x to the left, all other terms to the right. Add '20' to each side of the equation. -20 + 20 + 29x = 0 + 20 Combine like terms: -20 + 20 = 0 0 + 29x = 0 + 20 29x = 0 + 20 Combine like terms: 0 + 20 = 20 29x = 20 Divide each side by '29'. x = 0.6896551724 Simplifying x = 0.6896551724 Solution x = {0, 0.6896551724}

OpenStudy (anonymous):

x(29x-20)=0 x=0 >29x-20=0 >29x=20 >x=20/29 plug those back into the original equation and if it equals zero then it worked

OpenStudy (anonymous):

thank you guys, Now I have to post a few more questions. you wanna take a stab at them?

OpenStudy (anonymous):

@nathan917

OpenStudy (nathan917):

sure.